For each of the following hydroxide ion concentrations, what is the hydronium ion
concentration? The pH? The pOH? 3.5 × 10-5 M, 7.1 × 10-1 M, 4.5 × 10-10 M, 2.1 × 10-7
M
1
Expert's answer
2020-05-12T11:34:39-0400

"[H^+][OH^-]=10^{-14}"

"[H^+]=\frac{10^{-14}}{[OH^-]}"


pH=-log[H+]


1) [OH-] ="3.5\times10^{-5}"

[H+]="\frac{10^{-14




}}{3.5\times10^{-5}}"=2.85"\times"10-10


pH=-log[2.85"\times"10-10] = 10-0.45 =9.55